OpenStudy (anonymous):

7x^2-2=2 find the large root

OpenStudy (anonymous):

7x^2=4 sqrt(7x^2) = sqrt(4) 7x = 2 x = 2/7

OpenStudy (anonymous):

In first step, I took the -2 over to other side, so it became 4.

OpenStudy (mertsj):

\[7x^2=4\] \[x^2=\frac{4}{7}\] \[x=\pm\frac{2}{\sqrt{7}}=\pm\frac{2\sqrt{7}}{7}\]

OpenStudy (mertsj):

So we see that the large root is the positive one.

OpenStudy (anonymous):

so its 4 ?

OpenStudy (anonymous):

First lets move everything to one side, so that the right is 0: 7x^2 - 2 = 2 7x^2 - 2 -2 = 0 7x^2 - 4 = 0 Lets use the quadratic formula, since there is no bx, we know that b is 0: \[x = \sqrt{112} / 14 = 4\sqrt{7}/14 = 2\sqrt{7}/ 7\] so your answer is: \[2\sqrt{7}/7\]